# Towards the Usage of Window Counting Constraints

### v1

@@ 3 Incremental Synthesis with Counting Constraints
If the initial state of the adapted situation graph $Sit_{\mathcal{C}}$ is in the winning region, a set of winning strategies for $EGO$ is found for the game $G_{\mathcal{C}}=(Sit_{\mathcal{C}},\widetilde{\text{Win}})$. Any winning strategy in this game can easily be transformed to a (in general non-positional) winning strategy in $G=(A,\text{Win},CC)$. In this case, the incrementing process stops.
@@
4 Example and Experimental Results
@@
Following the process sketched in Figure 6, the first increment is using the counting constraint $CC_{min}(EGO,a,1,1)$ ($EGO$ plays $a$ at least in one of 1 turns”). In the initial state $1$, $EGO$ is only allowed to play $\neg a$, hence $EGO$ violates this constraint already with its first (and only) turn. Consequently, the only considered successor of the reached situation is the losing sink $L$ and the situation graph remains rather small, as depicted in Figure 9(a). Taking this situation graph as new game graph and winning condition $\widetilde{\text{Win}}=\mathtt{Safe(\{(1,(-)),(2,(0))\})}$ yields an empty winning region. For the next increment, the length of the counting constraint is increased by one, that is, $\{CC_{min}(EGO,a,1,2)\}$ is taken. The resulting situation graph is shown in Figure 9(b). Considering all of the included situations except for the two losing sink states as safe, the winning region of the resulting safety game now includes a subgraph with 10 situations and 12 transitions (marked in gray in the figure). Since the initial situation is not part of the winning region, no winning strategy for $EGO$ is found yet and the counting constraint is increased to $CC_{min}(EGO,a,1,3)$. The resulting situation graph is shown in 9(c). Note that this graph is rather small, since successors of situation extensions of the previous winning region do not need to be considered, but the extensions directly go into the winning sink. For example, the situation $(7,(0,-,-))$ is an extension of the situation $(7,(0,-))$, which is in the winning region of the second increment, hence the only successor of $(7,(0,-,-))$ is the winning sink. The same holds for situation $(6,(1,0,0))$ as extension of $(6,(1,0))$. With all situations in the third situation graph marked as safe (since no counting constraint violations are included), the whole graph becomes the winning region. Hence, the initial state is also winnable and a winning strategy for $EGO$ is found and no further increment (with potentially larger situation graphs) is needed.
@@
[…]
Figure 8: Two-player game graph $A$ for a safety game. States represented as circles are controlled by $EGO$, diamond-shaped states are controlled by $ALTER$. $\Sigma_{E}=\{a\}$ and outgoing transitions of states controlled by $EGO$ without label mean that $EGO$ plays $\neg a$ in those moves. The played letters of $ALTER$ are not important in this example, hence labels of outgoing transitions of $ALTER$-controlled states are omitted. To keep the example simple, we consider a safety game without unsafe states, that is, $\text{Win}=\mathtt{Safe(R)}$ for $R$ containing all states of $A$. $EGO$ shall fulfill the counting constraint $C=CC_{min}(EGO,a,1,7)$ ($EGO$ plays $a$ at least one time in 7 turns), making $G=(A,\text{Win},\{C\})$ a game with counting constraints.
@@
$\underset{(\text{-},\text{-})}{{1}}$$\underset{(0,\text{-})}{{2}}$$\underset{(0,\text{-})}{{3}}$$\underset{(0,\text{-})}{{7}}$$\underset{(0,0)}{{4}}$$\underset{(1,0)}{{8}}$$\underset{(1,0)}{{9}}$$\underset{(1,1)}{{8}}$$\underset{(1,1)}{{9}}$$\underset{(0,1)}{{10}}$$\underset{(0,1)}{{5}}$$\underset{(1,0)}{{6}}$$\underset{(1,0)}{{3}}$$\underset{(0,1)}{{4}}$$\underset{(\boldsymbol{\cdot})}{{l_{E}}}$$\underset{(\boldsymbol{\cdot})}{{l_{A}}}$$a$$a$$a$$a$
@@
Figure 9: Situation graphs for the game in Figure 8 with incrementing over $CC_{min}(EGO,a,1,7)$. Situations belonging to the winning region are marked in gray. Situations of the losing sink $L$ are marked double-framed and situations of the winning sink $W$ with dotted lines. Each of the showed situations (nodes in the graphs) consist of a state of the original game graph in Figure 8 and a history on how the situation was reached. For example, the history in situation $(6,(1,0,0))$ (upper right corner of Figure 9(c)) encodes that $EGO$ playes $a$ in its last move, but not in the second and third to last move. More than three increments are not necessary to wind a winning strategy for $EGO$, since there already is a winning strategy for $EGO$ in the third increment.
@@
Table 1: Runtime comparison of synthesis with and without incrementality
@@
No. Win? # $\mathcal{C}$ states time [sec] # states time [sec] # states time [sec]
@@ 5 Yes 2 537,970 111 152,882 51 ? ?
5 Discussion and Future Work
@@
[…]
Figure 10: $ALTER$ can always fulfill the counting constraint $CC_{min}(ALTER,b,1,3)$, but can run into a violation for $CC_{min}(ALTER,b,1,2)$

### v2

@@ 3. Incremental Synthesis with Counting Constraints
[…]
@@
4. Example and Experimental Results
@@
Following the process sketched in Figure 6, the first increment is using the counting constraint $CC_{min}(EGO,a,1,1)$ (EGOEGO plays $a$ at least in one of 1 turns”). In the initial state $1$, EGOEGO is only allowed to play $\neg a$, hence $EGO$ violates this constraint already with its first (and only) turn. Consequently, the only considered successor of the reached situation is the losing sink $L$ and the situation graph remains rather small, as depicted in Figure 9(a). Taking this situation graph as new game graph and winning condition $\widetilde{\text{Win}}=\mathtt{Safe(\{(1,(-)),(2,(0))\})}$ yields an empty winning region. For the next increment, the length of the counting constraint is increased by one, that is, $\{CC_{min}(EGO,a,1,2)\}$ is taken. The resulting situation graph is shown in Figure 9(b). Considering all of the included situations except for the two losing sink states as safe, the winning region of the resulting safety game now includes a subgraph with 10 situations and 12 transitions (marked in gray in the figure). Since the initial situation is not part of the winning region, no winning strategy for $EGO$ is found yet and the counting constraint is increased to $CC_{min}(EGO,a,1,3)$. The resulting situation graph is shown in 9(c). Note that this graph is rather small, since successors of situation extensions of the previous winning region do not need to be considered, but the extensions directly go into the winning sink. For example, the situation (7,(0,,))(7,(0,-,-)) is an extension of the situation (7,(0,))(7,(0,-)), which is in the winning region of the second increment, hence the only successor of (7,(0,,))(7,(0,-,-)) is the winning sink. The same holds for situation $(6,(1,0,0))$ as extension of $(6,(1,0))$. With all situations in the third situation graph marked as safe (since no counting constraint violations are included), the whole graph becomes the winning region. Hence, the initial state is also winnable and a winning strategy for $EGO$ is found and no further increment (with potentially larger situation graphs) is needed.
@@
1122334455667788991010aaaaaa
Figure 8. Two-player game graph $A$ for a safety game. States represented as circles are controlled by $EGO$, diamond-shaped states are controlled by $ALTER$. $\Sigma_{E}=\{a\}$ and outgoing transitions of states controlled by EGOEGO without label mean that $EGO$ plays $\neg a$ in those moves. The played letters of $ALTER$ are not important in this example, hence labels of outgoing transitions of ALTERALTER-controlled states are omitted. To keep the example simple, we consider a safety game without unsafe states, that is, $\text{Win}=\mathtt{Safe(R)}$ for $R$ containing all states of $A$. EGOEGO shall fulfill the counting constraint $C=CC_{min}(EGO,a,1,7)$ (EGOEGO plays $a$ at least one time in 7 turns), making $G=(A,\text{Win},\{C\})$ a game with counting constraints.
@@
$\underset{(\text{-},\text{-})}{{1}}$$\underset{(0,\text{-})}{{2}}$$\underset{(0,\text{-})}{{3}}$$\underset{(0,\text{-})}{{7}}$$\underset{(0,0)}{{4}}$$\underset{(1,0)}{{8}}$$\underset{(1,0)}{{9}}$$\underset{(1,1)}{{8}}$$\underset{(1,1)}{{9}}$$\underset{(0,1)}{{10}}$$\underset{(0,1)}{{5}}$$\underset{(1,0)}{{6}}$$\underset{(1,0)}{{3}}$$\underset{(0,1)}{{4}}$$\underset{(\boldsymbol{\cdot})}{{l_{E}}}$$\underset{(\boldsymbol{\cdot})}{{l_{A}}}$aaaaaaaa
@@
Figure 9. Situation graphs for the game in Figure 8 with incrementing over $CC_{min}(EGO,a,1,7)$. Situations belonging to the winning region are marked in gray. Situations of the losing sink $L$ are marked double-framed and situations of the winning sink $W$ with dotted lines. Each of the showed situations (nodes in the graphs) consist of a state of the original game graph in Figure 8 and a history on how the situation was reached. For example, the history in situation $(6,(1,0,0))$ (upper right corner of Figure 9(c)) encodes that EGOEGO playes $a$ in its last move, but not in the second and third to last move. More than three increments are not necessary to wind a winning strategy for $EGO$, since there already is a winning strategy for $EGO$ in the third increment.
@@
Table 1. Runtime comparison of synthesis with and without incrementality
@@
No. Win? # $\mathcal{C}$ # states time [sec] # states time [sec] # states time [sec]
@@ 5 Yes 2 537,970 111 152,882 51 ? ?
5. Discussion and Future Work
@@
Towards cooperative games: The presented approach did not consider incrementing over constraints of $ALTER$, but handled them as being fixed. The rationale for this lies in the property ALTERALTER cannot be forced into constraint violations” of games with counting constraints (Definition 2). In general, a game with counting constraints may satisfy the requirement of $ALTER$ always being able to adhere to its counting constraints, only to find the requirement violated for the game with a modified counting constraint as used in the increments. An illustrative example is provided in Figure 10. $ALTER$ has the counting constraint $CC_{min}(ALTER,b,1,3)$, i.e. $ALTER$ plays $b$ at least once in three of its turns. Recall that ALTERALTER cannot be forced into constraint violations” is defined as $ALTER$ is able to enlarge each prefix that satisfies the constraint such that the resulting prefix is also satisfying the constraint. This property is fulfilled when considering the game graph and the constraints $CC_{min}(ALTER,b,1,1)$ or $CC_{min}(ALTER,b,1,3)$. However, it is violated for $CC_{min}(ALTER,b,2,3)$, since the prefix $(1,a,2,\neg b,4,a,5)$ satisfies the constraint [^1] ALTERALTER could play $b$ forever to complete the prefix to an infinite play that satisfies the constraint. This play is not in $G$, but nonetheless is sufficient according to the definition of a prefix satisfying a constraint in Section 2., but there is no possibility for $ALTER$ to still satisfy the constraint with the next turn. Since the definition of a winning strategy relies on the game property of $ALTER$ not being forceable into counting constraint violations, 1 cannot be extended to increments over $ALTER$-constraints. However, such an extension would offer additional potential for more efficient synthesis algorithms.
@@
Figure 10. ALTERALTER can always fulfill the counting constraint $CC_{min}(ALTER,b,1,3)$, but can run into a violation for $CC_{min}(ALTER,b,1,2)$